ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 76 Debnath, Lokenath, University of Central Florida, U. S. A. Resonant Phenomena in Linear and Nonlinear Water Waves Some linear and nonlinear initial value problems in water of finite or infinite depth under the action of surface pressure moving at a constant speed will be presented. Attention will be given to evolution equations associated with these problems. A variational principle for Rossby waves in beta-plane approximation is formulated. Some conservation laws are discussed from full nonlinear equations of Rossby waves in oceans. References: Lokenath Debnath, Nonlinear Water Waves, Academic Press, Boston, 1995. NamChul Kim and Lokenath Debnath, Resonant wave interactions in a weakly viscous liquid, International Journal of Nonlinear Mechanics (to appear), 1998. Lokenath Debnath, Nonlinear Partial Differential Equations for Scientists and Engineers, Birkhauser, Verlag, Boston, 1997.